A firm produces two products, milk and cheese, Q1 and Q2 represent the output rate for milk and cheese respectively. The profit function is
π = -100+20Q1+60Q2-10Q1^2-2Q1Q2
Determine the rate of output for each product that will maximize profit.
1
Expert's answer
2015-08-19T00:00:44-0400
To determine the rate of output for each product that will maximize profit we need to find derivative from the profit function - marginal profit MP, which should be equal zero MP = 0.
MP(Q1) = 20 - 20Q1 - 2Q2 = 0
MP(Q2) = 60 - 2Q1 = 0
Q1 = 30
20 - 600 - 2Q2 = 0
Q2 = -290